OpenStudy (anonymous):

If f(x) = x^2-sqrt(x)+6, find f'x.

OpenStudy (anonymous):

\[f(x)=x ^{2} -\sqrt{x}+7\]

OpenStudy (anonymous):

have u tried it yet?

OpenStudy (anonymous):

I did this: \[f'(x) \lim_{h \rightarrow 0} \frac{ (a+h)^{2}-\sqrt{a+h}+7-a ^{2} +\sqrt{a} -7}{ h }\]

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

\[\frac{2ah+h ^{2} -\sqrt{a+h} +a}{ h }\] then, the professor told me to solve them separately. so I get:

OpenStudy (anonymous):

hey its: 2x -1/2 x ^-1/2 :)

OpenStudy (anonymous):

\[\frac{ 2a+h+h ^{2} }{ h } =2+h\]

OpenStudy (anonymous):

right! I just automatically write a's because thats the formula. ;)

OpenStudy (anonymous):

and for the other one is when I get a little bit confused. I just multiply by the conjugate. and get \[\frac{ 1}{ \sqrt{a+h} +\sqrt{a}}\] so if I combine them I get: \[2+h -\frac{ 1 }{\sqrt{a+h} +\sqrt{a} }\]

OpenStudy (anonymous):

is this correct??